Researchers are often interested in the rate of political participation based on neighborhood economic conditions. Create a scatter plot of the variables and estimate the correlation between these tw...

  1. Researchers are often interested in the rate of political participation based on neighborhood economic conditions. Create a scatter plot of the variables and estimate the correlation between these two variables (assume each is continuous).

Neighborhood

Economic Conditions

Political Participation

1

36.3

16.4

2

29.4

14.8

3

45.6

68.7

4

65.3

78.8

5

66.8

77.4

6

98.4

65.1

7

41.3

55.4

8

12.6

9.6

9

9.5

22.1

10

89.6

98.6

11

62.1

45.6

12

45.6

37.3

13

68.5

56.4

14

72.5

64.1

15

39.8

25.7

16

25.7

17.5

17

61.4

72.1

18

55.6

61.8

19

44.8

48.8

20

48.9

51.5

21

51.3

60.1

22

55.3

62.7
Answer #1

Let X : Economic Conditions

Y: Political Participation

Following is the scatter plot of the data:

120 100 80 60 40 20 0 100 120 20 40 0 60 80

Following table shows the calculations for correlation coefficient:

X Y X^2 Y^2 XY
36.3 16.4 1317.69 268.96 595.32
29.4 14.8 864.36 219.04 435.12
45.6 68.7 2079.36 4719.69 3132.72
65.3 78.8 4264.09 6209.44 5145.64
66.8 77.4 4462.24 5990.76 5170.32
98.4 65.1 9682.56 4238.01 6405.84
41.3 55.4 1705.69 3069.16 2288.02
12.6 9.6 158.76 92.16 120.96
9.5 22.1 90.25 488.41 209.95
89.6 98.6 8028.16 9721.96 8834.56
62.1 45.6 3856.41 2079.36 2831.76
45.6 37.3 2079.36 1391.29 1700.88
68.5 56.4 4692.25 3180.96 3863.4
72.5 64.1 5256.25 4108.81 4647.25
39.8 25.7 1584.04 660.49 1022.86
25.7 17.5 660.49 306.25 449.75
61.4 72.1 3769.96 5198.41 4426.94
55.6 61.8 3091.36 3819.24 3436.08
44.8 48.8 2007.04 2381.44 2186.24
48.9 51.5 2391.21 2652.25 2518.35
51.3 60.1 2631.69 3612.01 3083.13
55.3 62.7 3058.09 3931.29 3467.31
Total 1126.3 1110.5 67731.31 68339.39 65972.4

Sample size: n=22

Now,

S_{yy}=\sum y^{2}-\frac{\left (\sum y \right )^{2}}{n}=12284.37864

S_{xx}=\sum x^{2}-\frac{\left (\sum x \right )^{2}}{n}=10069.86955

S_{xy}=\sum xy-\frac{\left (\sum x \right )\left (\sum y \right )}{n}=9119.847727

The coefficient of correlation is :

r=\frac{S_{xy}}{\sqrt{S_{xx}S_{yy}}}=0.82

Scatter plot shows that there is a strong linear relationship between the variables and correlation coefficient also shows the same.